# Solution to a Problem of Erdős Concerning Distances and Points

### v1

@@ Abstract.
In 1997, Erdős asked [1] whether for arbitrarily large $n$ there exists a set of $n$ points in $\mathbb{R}^{2}$ that determines $O(\frac{n}{\sqrt{\log n}})$ distinct distances while satisfying the local constraint that every 4-point subset determines at least 3 distinct pairwise distances. We construct $n$-point sets from an $m\times m$ box of the lattice
@@
The distinct distance bound follows from applying Bernays’ theorem [3] to the number of integers represented by the binary quadratic form $u^{2}+2v^{2}$. The local 4-point constraint is verified through Perucca’s similarity classification [5] of the six similarity types determining exactly two distances.
@@ 1. Introduction
Our approach uses a standard number-theoretic argument, applying Bernays’ theorem on integers represented by binary quadratic forms to a representation of the lattice’s distinct distances. While the distance counting method is classical, the present note demonstrates that a specific anisotropic variant achieves this global bound while preserving the local constraint. We verify this by a case analysis of the six configurations given by Perucca [5]. Our simple construction proves the following theorem, answering Erdős’ question in the affirmative.
@@ 5. Conclusion
and the local constraint that every 44-point subset determines at least three distinct pairwise distances. For a general $n$, taking any $n$-point subset of $P_{\lceil\sqrt{n}\rceil}$ preserves the local constraint and gives the same asymptotic upper bound (up to constants), as in Corollary 4. The construction ultimately works because $60^{\circ}$ and $90^{\circ}$ rotations, as well as the golden-ratio distance relation, are incompatable with with the quadratic field $\mathbb{Q}(\sqrt{2})$.
@@ Acknowledgements
Gemini 3.0 provided the core ideation behind this approach. The author relied extensively on GPT-5.2 and Gemini 3.0 for drafting, editing, and exploring possible proof strategies. All mathematical claims, proofs, and citations were independently verified by the author, who takes responsibility for any errors. I am grateful to Professor Peter Doyle and Professor Peter Winkler for helpful discussions and feedback on early drafts.

### v2

@@ Abstract.
In 1997, Erdős asked [1] whether for arbitrarily large $n$ there exists a set of $n$ points in $\mathbb{R}^{2}$ that determines $O(\frac{n}{\sqrt{\log n}})$ distinct distances while satisfying the local constraint that every 44-point subset determines at least 33 distinct pairwise distances. We use $n$-point subsets of an $m\times m$ box in the anisotropic lattice
@@
The distinct-distance bound follows from Bernays’ theorem applied to the binary quadratic form $u^{2}+2v^{2}$ [3], following earlier work on lattices with few distances [5, 6]. The local 44-point constraint is verified using Perucca’s classification [7] of 44-point two-distance sets, together with the observation noted here that this lattice excludes the remaining regular-pentagon trapezoid configuration.
@@ 1. Introduction
Constructions of planar point sets with few distinct distances arising from anisotropic lattices, together with the use of Bernays’ theorem (or related results) to count represented integers, appear in earlier work of Moree and Osburn [6]; see also Sheffer’s discussion [5]. The additional point in the present note is that the specific lattice $\mathbb{Z}\times\sqrt{2}\,\mathbb{Z}$ also excludes the remaining 44-point two-distance configuration consisting of four vertices of a regular pentagon (equivalently, the associated isosceles trapezoid). Combined with the complete similarity classification of 44-point two-distance sets [7], this verifies the local constraint, answering Erdős’ question in the affirmative.
@@ 4.2. Equilateral triangles
Remark.
The exclusion of squares and equilateral triangles in rectangular or anisotropic lattice constructions is standard; see [5] for related discussion.
@@ 5. Conclusion
and the local constraint that every 44-point subset determines at least three distinct pairwise distances. For a general $n$, taking any $n$-point subset of $P_{\lceil\sqrt{n}\rceil}$ preserves the local constraint and gives the same asymptotic upper bound (up to constants), as in Corollary 4. The construction ultimately works because $60^{\circ}$ and $90^{\circ}$ rotations, as well as the golden-ratio distance relation, are incompatible with the quadratic field $\mathbb{Q}(\sqrt{2})$ and this last obstruction rules out the regular-pentagon trapezoid configuration.
@@ Acknowledgements
Gemini 3.0 provided the core ideation behind this approach, and specifically surfaced the observation that the same lattice excludes the regular-pentagon trapezoid configuration. The author relied extensively on GPT-5.2 and Gemini 3.0 for drafting, editing, and exploring possible proof strategies. All mathematical claims, proofs, and citations were independently verified by the author, who takes responsibility for any errors. I am grateful to Professor Peter Doyle and Professor Peter Winkler for helpful discussions and feedback on early drafts. I also thank the participants in the public comment thread on the Erdős Problems website [2], particularly Desmond Weisenberg for identifying the local constraint as six forbidden shapes, Terrence Tao for highlighting prior work from Sheffer and Moree–Osburn, and Boris Alexeev for formalizing the proof in Lean.